# Beyond In-Domain Detection: SpikeScore for Cross-Domain Hallucination Detection

### v1

@@ 2 Learning Setups
Generalizable Hallucination Detection. In GHD, we consider a training domain $\mathbb{P}_{Q,T}$ and $N$ related test domains $\mathbb{P}_{Q,T}^{1},\ldots,\mathbb{P}_{Q,T}^{N}$. The goal of GHD is to train a detector on $\mathbb{P}_{Q,T}$ that generalizes well to the $N$ test domains. We formally introduce the definition of GHD below.
@@
as introduced in Eq. (2), the objective of generalizable hallucination detection (GHD) is to learn a detector $D$ based on the LLM $\mathbb{P}_{\bm{\theta}}(\cdot)$ and the training dataset $\mathcal{D}_{l}$. The detector is required to generalize across $N$ test domains $\mathbb{P}_{Q,T}^{1},\ldots,\mathbb{P}_{Q,T}^{N}$, i.e., for any question–answer pair $(\mathbf{Q},\mathbf{A})$,
@@ 3.1 Exploring Cross-Domain Invariant Indicators
$\mathbf{S}(\mathbf{Q},\mathbf{A})=[S(\mathbf{A}^{1};\mathbf{Q},\mathbf{A}),S(\mathbf{A}^{2};\mathbf{Q},\mathbf{A}),...,S^{K}(\mathbf{A}^{1};\mathbf{Q},\mathbf{A})],$ (8)
@@ 3.2 Evaluations of SpikeScore
Let $\mathbb{P}_{Q,T}^{t}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{P}_{Q,T}^{i}$ be the non-hallucination domain as the uniform mixture of $N$ test domains $\mathbb{P}_{Q,T}^{1},...,\mathbb{P}_{Q,T}^{N}$ introduced in Section 2. The value of SpikeScore trained on the associated training domain $\mathbb{P}_{Q,T}$ satisfies that
@@
In Figure 2 (b), we observe that the standard deviation of SpikeScore in the hallucination domain is larger than in the non-hallucination domain, with the ratio of standard deviations reaching $2.5$, which exceeds the expectation ratio of $2$ reported in Observation 1. This indicates that variance differences alone may hinder direct separability. Nevertheless, as shown in Theorem 1, separability between hallucination and non-hallucination domains can still be established under controlled coefficient-of-variation conditions (i.e., variance normalized by expectation). See Theorem 1 for details.
@@ 3.3 Leveraging SpikeScore for GHD
where $1$ indicates that $\mathbf{A}$ is hallucinated, and 0 indicates that $\mathbf{A}$ is non-hallucinated.
@@ 4.1 Experimental Setups
We evaluate SpikeScore on six widely used hallucination detection benchmarks: CommonsenseQA (talmor-etal-2019-commonsenseqa), TriviaQA (joshi-etal-2017-triviaqa), Belebele (costa-jussa-etal-2025-2m), CoQA (reddy-etal-2019-coqa), Math (hendrycks2021measuring), and SVAMP (patel-etal-2021-nlp). Baselines include Perplexity (ren2023outofdistribution), Semantic Entropy (farquhar2024detecting), EigenScore (chen2024inside), Lexical Similarity (lin2024generating), Verbalize (lin2022teaching), InterrogateLLM (yehuda-etal-2024-interrogatellm), MM (marks2023geometry), SEP (kossen2024semantic; farquhar2024detecting), SAPLMA (azaria2023internal), PRISM (SAPLMA instantiation) (zhang2024prompt), and ICR Probe (zhang-etal-2025-icr), with AUROC as the primary evaluation metric. We test across four carefully selected open-source LLMs—Llama-3.2-3B-Instruct and Llama-3.1-8B-Instruct (abs-2407-21783), and Qwen3-8B-Instruct and Qwen3-14B-Instruct (qwen3)—to cover diverse model families and parameter scales. Additional dataset sampling strategies, baseline configurations, and hyperparameters are provided in detail in Appendix LABEL:app:details.
@@ 4.2 Experimental Results
Overall Cross-domain Hallucination Detection Performance. In Table 3.3, we compare the proposed SpikeScore with three categories of baselines: training-free methods that avoid generalization issues, traditional training-based methods, and cross-domain specialized approaches. SpikeScore consistently achieves the highest average AUROC across all model families and maintains superior or competitive performance in individual dataset evaluations. Traditional training-based methods (MM, SEP, SAPLMA) generally exhibit poor cross-domain performance, while training-free methods show reasonable adaptability but fall short of specialized cross-domain approaches like PRISM and ICR Probe. Notably, SpikeScore demonstrates improved performance with larger models within the same family, suggesting that enhanced self-correction capabilities in larger LLMs (yang-etal-2025-confidence) generate more detectable attention spike patterns during conversational refinement.
@@ Appendix E Discussion on Separability
    Formally, for a single domain $A$ with hallucination and non-hallucination distributions $(X_{A},Y_{A})$, Cantelli’s inequality yields
@@
    where $\delta_{A}=\tfrac{\mathbb{E}X_{A}}{\mathbb{E}Y_{A}}$ is the mean ratio, $r_{A}=\tfrac{\mathrm{Std}(X_{A})}{\mathrm{Std}(Y_{A})}$ is the variance ratio, and $c_{A}=\tfrac{\mathrm{Std}(Y_{A})}{\mathbb{E}Y_{A}}$ is the coefficient of variation of non-hallucinations. If domain $A$ happens to exhibit weaker mean gaps ($\delta_{A}\!\approx\!1$), larger variance inflation ($r_{A}$ large), or higher non-hallucination noise ($c_{A}$ large), then its AUROC can drop below the global mixture bound. In practice, we indeed observe in Table 3.3 and Table E that datasets such as Belebele occasionally yield lower per-domain AUROCs relative to the theoretical bound. This does not contradict our analysis; rather, it reflects the fact that separability is guaranteed only at the mixture level, and domains with less favorable statistical profiles naturally appear weaker. Intuitively, these domains are “harder” because hallucinations and factual answers are less distinguishable in terms of score dynamics. Hence, the occasional $\mathrm{AUROC}_{A}<\mathrm{LB}_{\mathrm{mix}}$ is a theoretically anticipated phenomenon and does not necessitate further ablations.

### v2

@@ 2 Learning Setups
Generalizable Hallucination Detection. In GHD, we consider a training domain $\mathbb{P}_{Q,T}$ and NN related test domains $\mathbb{P}_{Q,T}^{1},\ldots,\mathbb{P}_{Q,T}^{N}$. The goal of GHD is to train a detector on $\mathbb{P}_{Q,T}$ that generalizes well to the NN test domains. We formally introduce the definition of GHD below.
@@
as introduced in Eq. (2), the objective of generalizable hallucination detection (GHD) is to learn a detector $D$ based on the LLM $\mathbb{P}_{\bm{\theta}}(\cdot)$ and the training dataset $\mathcal{D}_{l}$. The detector is required to generalize across NN test domains $\mathbb{P}_{Q,T}^{1},\ldots,\mathbb{P}_{Q,T}^{N}$, i.e., for any question–answer pair $(\mathbf{Q},\mathbf{A})$,
@@ 3.1 Exploring Cross-Domain Invariant Indicators
$\mathbf{S}(\mathbf{Q},\mathbf{A})=\big[S(\mathbf{A}^{1};\mathbf{Q},\mathbf{A}),\,S(\mathbf{A}^{2};\mathbf{Q},\mathbf{A}),\,\ldots,\,S(\mathbf{A}^{K};\mathbf{Q},\mathbf{A})\big].$ (8)
@@ 3.2 Evaluations of SpikeScore
Let $\mathbb{P}_{Q,T}^{t}=\frac{1}{N}\sum_{i=1}^{N}\mathbb{P}_{Q,T}^{i}$ be the non-hallucination domain as the uniform mixture of NN test domains $\mathbb{P}_{Q,T}^{1},...,\mathbb{P}_{Q,T}^{N}$ introduced in Section 2. The value of SpikeScore trained on the associated training domain $\mathbb{P}_{Q,T}$ satisfies that
@@
In Figure 2 (b), we observe that the standard deviation of SpikeScore in the hallucination domain is larger than in the non-hallucination domain, with the ratio of standard deviations reaching $2.5$, which exceeds the expectation ratio of 22 reported in Observation 1. This indicates that variance differences alone may hinder direct separability. Nevertheless, as shown in Theorem 1, separability between hallucination and non-hallucination domains can still be established under controlled coefficient-of-variation conditions (i.e., variance normalized by expectation). See Theorem 1 for details.
@@ 3.3 Leveraging SpikeScore for GHD
where 11 indicates that $\mathbf{A}$ is hallucinated, and 0 indicates that $\mathbf{A}$ is non-hallucinated.
@@ 4.1 Experimental Setups
We evaluate SpikeScore on six widely used hallucination detection benchmarks: CommonsenseQA (talmor-etal-2019-commonsenseqa), TriviaQA (joshi-etal-2017-triviaqa), Belebele (costa-jussa-etal-2025-2m), CoQA (reddy-etal-2019-coqa), Math (hendrycks2021measuring), and SVAMP (patel-etal-2021-nlp). Baselines include Perplexity (ren2023outofdistribution), Semantic Entropy (farquhar2024detecting), EigenScore (chen2024inside), Lexical Similarity (lin2024generating), Verbalize (lin2022teaching), InterrogateLLM (yehuda-etal-2024-interrogatellm), $M$ (marks2023geometry), SEP (kossen2024semantic; farquhar2024detecting), SAPLMA (azaria2023internal), PRISM (SAPLMA instantiation) (zhang2024prompt), and ICR Probe (zhang-etal-2025-icr), with AUROC as the primary evaluation metric. We test across four carefully selected open-source LLMs—Llama-3.2-3B-Instruct and Llama-3.1-8B-Instruct (abs-2407-21783), and Qwen3-8B-Instruct and Qwen3-14B-Instruct (qwen3)—to cover diverse model families and parameter scales. Additional dataset sampling strategies, baseline configurations, and hyperparameters are provided in detail in Appendix LABEL:app:details.
@@ 4.2 Experimental Results
Overall Cross-domain Hallucination Detection Performance. In Table 3.3, we compare the proposed SpikeScore with three categories of baselines: training-free methods that avoid generalization issues, traditional training-based methods, and cross-domain specialized approaches. SpikeScore consistently achieves the highest average AUROC across all model families and maintains superior or competitive performance in individual dataset evaluations. Traditional training-based methods ($M$, SEP, SAPLMA) generally exhibit poor cross-domain performance, while training-free methods show reasonable adaptability but fall short of specialized cross-domain approaches like PRISM and ICR Probe. Notably, SpikeScore demonstrates improved performance with larger models within the same family, suggesting that enhanced self-correction capabilities in larger LLMs (yang-etal-2025-confidence) generate more detectable attention spike patterns during conversational refinement.
@@ Appendix E Discussion on Separability
    Formally, for a single domain AA with hallucination and non-hallucination distributions $(X_{A},Y_{A})$, Cantelli’s inequality yields
@@
    where $\delta_{A}=\tfrac{\mathbb{E}X_{A}}{\mathbb{E}Y_{A}}$ is the mean ratio, $r_{A}=\tfrac{\mathrm{Std}(X_{A})}{\mathrm{Std}(Y_{A})}$ is the variance ratio, and $c_{A}=\tfrac{\mathrm{Std}(Y_{A})}{\mathbb{E}Y_{A}}$ is the coefficient of variation of non-hallucinations. If domain AA happens to exhibit weaker mean gaps ($\delta_{A}\!\approx\!1$), larger variance inflation ($r_{A}$ large), or higher non-hallucination noise ($c_{A}$ large), then its AUROC can drop below the global mixture bound. In practice, we indeed observe in Table 3.3 and Table E that datasets such as Belebele occasionally yield lower per-domain AUROCs relative to the theoretical bound. This does not contradict our analysis; rather, it reflects the fact that separability is guaranteed only at the mixture level, and domains with less favorable statistical profiles naturally appear weaker. Intuitively, these domains are “harder” because hallucinations and factual answers are less distinguishable in terms of score dynamics. Hence, the occasional $\mathrm{AUROC}_{A}<\mathrm{LB}_{\mathrm{mix}}$ is a theoretically anticipated phenomenon and does not necessitate further ablations.